ques) A thief in a stolen car through a police check post at his top speed of 90 km/h. A motorcycle , reacting after 2 sec, accelerates from rest at 5 m/s2. His top speed being 108 km/h. Find the max seperation between policeman and thief.


 

Grade:12

2 Answers

Prasanjeet Mohanty Mohanty
14 years ago

After 2s the thief would be at a distance of 50m fromthe police car.We shall consider this to be t=0.

let after a time t the distance be max.

then the distance between the check post and the thief would be 50+25t.......(i)

the police car would be at a distance of 2.5t^2from the check post...........(ii)

 then the distance between them would be x=50+25t-2.5t^2

this would be max when dx/dt=0

which implies dx/dt=25-5t=0

therefore t=5s.hence x=50+125-62.5=112.5(ans)

Prasanjeet Mohanty Mohanty
14 years ago

P.s I have converted the velocity from Km/h to m/s